OpenStudy (anonymous):

fine b and c so that y=7x^2+bx+c has a vertex (4,-10)

OpenStudy (campbell_st):

using the line of symmetry, b = 4 and using the equation for the line of symmetry you get \[4 = \frac{-b}{2 \times 7}\] just solve for b when you find b, put it into the equation. then substitute the vertex, x = 4 and y = -10 into the equation to find c hope it helps

OpenStudy (campbell_st):

well I got b = -56 so you have \[-10 = 7 \times 4^2 - 56 \times 4 + c \] that should help you find c